Solve for the values of x for each of the following exponential equations and inequalities.


1. 82-2


2. () <=


3. 5* = 25-2


4. 3x+2 à ¢ ‰ ¥ 27


5. 43x 8-1


1
Expert's answer
2021-10-18T15:08:56-0400

1.

"8^{2-x}=2"

"8^{2-x}=8^{1\/3}"

"2-x=\dfrac{1}{3}"

"x=\dfrac{5}{3}"

2.

"(\dfrac{1}{2})^{x}<\dfrac{1}{8}"

"(\dfrac{1}{2})^{x}<(\dfrac{1}{2})^{3}"

"x>3"

3.

"5^x=25^{x-2}"

"25^{x\/2}=25^{x-2}"

"\dfrac{x}{2}=x-2"

"\dfrac{x}{2}=2"

"x=4"


4.

"3^{x+2}\geq27"

"3^{x+2}\geq3^3"

"x+2\geq3"

"x\geq1"

5.

"4^{3x}=8^{x-1}"

"2^{6x}=2^{3x-3}"

"6x=3x-3"

"3x=-3"

"x=-1"